OpenStudy (tori040403):

Please help will fan and medal! Which shows three ordered pairs that are all solutions of the equation y = x + 7? A. (0, 7), (2, 9), and (3, 11) B. (0, 7), (2, 9), and (3, 10) C. (1, 7), (2, 9), and (3, 10) D. (1, 8), (2, 9), and (3, 11)

OpenStudy (tori040403):

@AnswerMyQuestions

OpenStudy (tori040403):

@KrazyKricket

OpenStudy (anonymous):

Not sure... :/

OpenStudy (tori040403):

:/

OpenStudy (tori040403):

im listening to this day aria from mlp XD

OpenStudy (texaschic101):

y = x + 7 now we start subbing (0,7)...x = 0 and y = 7 7 = 0 + 7 7 = 7 (correct) (2,9)...x = 2 and y = 9 9 = 2 + 7 9 = 9 (correct) (3,11)...x = 3 and y = 11 11 = 3 + 7 11 = 10 (wrong) so we know it is not A because (3,11) is not a solution...but (2,9) and (0,7) are.. Do you understand how I am doing this ? I am just subbing in my points to see which one satisfies the equation.